What is the probability of extinction for a species of rare bird with an initial population size of 118, b = 1.39, d = 1.31. What if there were only 20 birds to start?
Answer #1

Initial population size N0 = 118 or N0 = 20.

Birth rate b = 1.39, death rate d = 1.31 and b > d.

The probability of extinction E can be calculated from

  PE = \left( \frac{d}{b}\right )^{N_0}

When N0 = 118, PE = (1.31/1.39)118 = 9.17 x 10-4.

When N0 = 20, PE = (1.31/1.39)20 = 0.306.